Question 1 (1 point)

Suppose the consumer's utility function is given by

U(x1,x2)=x1ax2b


, where a=7.0, and b=4.0. Suppose further, the price of good 1 is 2.6, price of good 2 is 0.5 and income is 43

Finally suppose the price of good 1 has decreased to 0.4

What is the substitution effect for good 1 and good 2? Round your answer to two decimal places

Suppose the consumer's utility function is given by

U(x1,x2)=x1ax2b


, where a=2.0, and b=5.5. Suppose further, the price of good 1 is 1.4, price of good 2 is 2.7 and income is 94

Finally suppose the price of good 1 has decreased to 0.4

What is the income effect for good 1 and good 2 Round your answer to two decimal places
Answer #1

Here the consumer utility fr. is. (x, Rr) - su ab where az 7.0 and b=4.0. Alo Given price of good 1 (b) = 2.6 price of good 2

As p. = 2.6 $2=0. 5 m = 43. 2.6 x + 0.5 X2 = 43. when &q=0. 2.6% = 43 x =43 = 16.54 26 when Noo 0.5 X2=43 d2 = 43 - 86 Os RefNow the price of good I decreases to 0.4 but the income remain the same so the real income of the consumer increases. So to f

Now the equation of budget line is alle b, x, it ba az a As p = 0.4 p 2 = 0.5 = 31.57 so 0.484 t o. 582= 31.57 eur of the con

3.7 = 220.99 0.5 5.5 - 40.18 good I falls. So Previously optimal bundle for was 10.52 but as price of good i optimal bundle i